cigarettes are forbidden, so people trade cigarettes in a black market. TheΒ 

cigarette demand is 𝑄𝐷 = 12 βˆ’ 𝑝, and cigarette supply is 𝑄𝑆 = 2𝑝.


1
Expert's answer
2021-05-31T15:00:12-0400

From the given information, equilibrium price and equilibrium quantity can be calculated.


At equilibrium, "Q_d=Q_s"

"12-P=2P"

"3P=12"

"equilibrium \space price P=4"

"equilibrium\space quantity\space Q=12-4=8"